OpenStudy (anonymous):

Can you check if I am correct? 10^2 ÷ 25 • 20 + 4 ÷ 2 – 14 = 82

Mehek (mehek14):

did you use PEMDAS?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

forgot something

OpenStudy (anonymous):

68

Mehek (mehek14):

you have 68 as your final answer?

OpenStudy (anonymous):

yup since 100/25 = 4 and 4 * 20 = 80 and 4/2 = 2 and 80 + 2 = 82 and 82 - 14 = 68

OpenStudy (anonymous):

I just forgot the last part of it

Mehek (mehek14):

68 is correct :)

OpenStudy (anonymous):

thx just wanted to clarify

OpenStudy (anonymous):

5 • [(14 + 3 • 30) + 120] + 103 – 10

OpenStudy (anonymous):

= 2110

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

btw 103 = 10^3

OpenStudy (anonymous):

sorry for it being confusing confusing

Mehek (mehek14):

it's correct too :)

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw

OpenStudy (anonymous):

how about 45 + 22 ÷ 11 • {[(60 – 15) • 60] + 25} could you help me solve this?

Mehek (mehek14):

60 - 15 first

OpenStudy (anonymous):

kk since it is inner

OpenStudy (anonymous):

the inner one

Mehek (mehek14):

ye

Mehek (mehek14):

then * 60